OpenStudy (anonymous):

Find the length of the longer diagonal of the rhombus with sides of 6cm and acute angles of 40 degrees.

OpenStudy (anonymous):

nvrmind

OpenStudy (anonymous):

Umm ok.

OpenStudy (anonymous):

I guess what I have to do is divide the rhombus into two triangles. Then what do I do?

OpenStudy (anonymous):

And I guess the hypotenuse of one these triangles would be the diaganol I'm looking for correct?

OpenStudy (anonymous):



OpenStudy (anonymous):

the length of the longer diagonal is 2y. since the diagonals of a rhombus are perpendicular and bisect each other we can just focus on 1 triangle. notice that cos(20 degrees) = y/6. so y=6*cos(20 degrees). the length of the longer diagonal is therefore 2y = 2*6cos(20 degrees) = 12*cos(20 degrees)

OpenStudy (anonymous):

All in all it was just a brick in the wall...

OpenStudy (anonymous):

Haha.

OpenStudy (anonymous):

Woah man thanks. That helps. For some reason I thought it would be like this.